For this problem I need to complete the square, and find the roots of the quadratic equation
x^2 + 16x = 0

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

complete the square, and find the roots of the quadratic equation
x^2 + 16x = 0
---
x^2 + 16x + 64 = 64
(x+8)^2 = 64
x+8 = 8 or x+8 = -8
x = 0 or x = -16
---------------------
Cheers,
Stan H.

Answer by solver91311(24713)



Step 1: Put the equation into standard form, that is: . (This is already done for this problem)

Step 2: Multiply both sides of the equation by the reciprocal of the lead coefficient, that is: . becomes . Since the lead coefficient in the given problem is 1, this step can be skipped.

Step 3: Add the additive inverse of the constant term to both sides, that is: becomes . (This is already done for this problem)

Step 4: Divide the coefficient on the 1st degree term by 2 and then square the result, that is, calculate . For this problem: 16 divided by 2 is 8, 8 squared is 64.

Step 5: Add the results of step 3 to both sides of the equation, that is: becomes For this problem, .

Step 6: Factor the perfect square on the left, that is: becomes . For this problem: .

Step 7: Take the square root of both sides. Remember to consider both the positive and negative roots, that is: becomes . For this problem:

Step 8: Collect terms and simplify, that is: becomes . For this problem: , so or .

John